OpenStudy (anonymous):

Write an equation. When 2 less than 3 lots of a number is doubled the result is 5.

OpenStudy (anonymous):

i think it should be \[(2-3x)^{2}=5\]

OpenStudy (anonymous):

that would mean x = -1/6?

OpenStudy (anonymous):

dont worry i found the answer, the equation is 2(3x-2)=5

OpenStudy (anonymous):

okay then.....goodluck!......:)